OpenStudy (anonymous):

The base of a rectangular prism is 5.4 centimeters long and 4.9 centimeters wide. It is 3.8 centimeters tall. What is the volume of the prism? Round your answer to the nearest tenth. Select the best answer from the choices provided. 100.5 cm3 110.8 cm3 129.7 cm3 77.9 cm3

OpenStudy (alexandervonhumboldt2):

multiply l**w*t=5.4*4.9*3.8=????

OpenStudy (anonymous):

100.549? @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

no 100.5